OpenStudy (anonymous):

Plz help me understand. The lesson didn't make sense The lengths of two sides of a triangle are 12 inches and 4 inches. Which of the following dimensions is the third side of this triangle? 9 inches 8 inches 7 inches 6 inches

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Third scale means the length of the missing side of the triangle

OpenStudy (anonymous):

in a triangle, sum of any two sides is greater than the third side  x+y>z x+z>y y+z>x

OpenStudy (anonymous):

sooo...yeah I am still confused. I don't want the answer directly but can u plz use the numbers given lolol

OpenStudy (anonymous):

ok , so we already know that the biggest side is 12( because all the options given are also less than 12) so the third side should be such that , when added to 4, gives a number bigger than 12

OpenStudy (anonymous):

And 9 is the only measurement that meets that condition.

OpenStudy (anonymous):

uh. u just gave the answer but yet my slow brain is still confused as to why thats the answer XDXD

OpenStudy (anonymous):

the sum of any two sides should be bigger than the third side if the third side is 6, 6+4=10 , which is not bigger than 12 similarly 4+7=11 4+8=12 which are not bigger than 12 only when you add 9+4=13 gives a number bigger than 12

OpenStudy (anonymous):

OMG I GET IT THANK YOU THANK YOU THANK YOU

OpenStudy (anonymous):

7 inches

OpenStudy (anonymous):

u r welcome ! :)